OpenStudy (anonymous):

You place a balloon in a closed chamber at standard temperature and pressure (STP). You double the chamber temperature while leaving the pressure constant. What happens to the balloon? It doubles in volume. It decreases to one-half of its original volume. It stays the same volume. It expands to four times its original volume.

OpenStudy (anonymous):

@YOUNGSUPERMAN

OpenStudy (youngsuperman):

b

OpenStudy (anonymous):

you sure? And can you help with more? I have a lot ;-;

OpenStudy (anonymous):

@Baby_bear

OpenStudy (anonymous):

@baby_bear69

OpenStudy (anonymous):

You place air in a sealed can at standard temperature and pressure (STP). You triple its absolute temperature (K) while leaving the volume constant. What happens to the pressure? (Points : 3) It triples. It decreases by one-half. It stays the same. It decreases to one-third of the original pressure.

OpenStudy (youngsuperman):

u still need a little help

OpenStudy (anonymous):

Now Im confused @just_one_last_goodbye

OpenStudy (youngsuperman):

@Baby_Bear69 did u get wat i said

OpenStudy (baby_bear69):

Mike is right, the first one is B

OpenStudy (anonymous):

Okay, this is so confusing >_<

OpenStudy (baby_bear69):

the second one is A

OpenStudy (anonymous):

you sure?

OpenStudy (just_one_last_goodbye):

o.o hey babe which one u want me to do? (for that very special reward better than a medal)

OpenStudy (anonymous):

Welp you wont get it for just one question ;)

OpenStudy (anonymous):

Which of the following is the mathematical representation of Boyle's law? (Points : 3) P1V1 = P2V2 P1V1 = V2/T2 V1/T1 = V2/T2 P1/T1 = P2/T2

OpenStudy (baby_bear69):

o.o

OpenStudy (just_one_last_goodbye):

o.o gotta be kidding me bruh...

OpenStudy (just_one_last_goodbye):

got a probelm with something @Baby_Bear69

OpenStudy (baby_bear69):

no?

OpenStudy (anonymous):

Nope. Im not.

OpenStudy (just_one_last_goodbye):

good =.=

OpenStudy (baby_bear69):

we used to be friends.. what happened.? & Leila, the first two that i answered i'm pretty sure are right..

OpenStudy (anonymous):

Okay and the third? Im sorry angel but when it comes to chem im a dimwit >_>

OpenStudy (just_one_last_goodbye):

Third is P1V1 = p2V2 (choice A)

OpenStudy (anonymous):

Okay, thank you!! You guys up for helping me finish this? At room temperature, which of the following is the most compressible? (Points : 3) one liter of mercury, a liquid one liter of iron, a solid one liter of oxygen, a gas one liter of water, a liquid

OpenStudy (anonymous):

its 15 problems o-o

OpenStudy (youngsuperman):

c i think

OpenStudy (just_one_last_goodbye):

^ agreed Only gasses are compressible.

OpenStudy (anonymous):

okay thank youuuu!!! You open a can of soda at room temperature and hear a hiss. Which of the following factors has changed inside the container? (Points : 3) pressure of gas temperature of gas type of gas atmospheric pressure

OpenStudy (anonymous):

A?

OpenStudy (baby_bear69):

i would say A

OpenStudy (just_one_last_goodbye):

I would say D

OpenStudy (anonymous):

...o-o?

OpenStudy (youngsuperman):

yea meetoo

OpenStudy (anonymous):

So its D?

OpenStudy (just_one_last_goodbye):

hey is did i reach my reward yet? Leila

OpenStudy (baby_bear69):

I'm sticking with A but D might be right

OpenStudy (anonymous):

Soon captain...soon xD I think D makes a little more sense

OpenStudy (somy):

@LeilaJudeh One question per post please!

OpenStudy (anonymous):

=_=

OpenStudy (somy):

+ second question you got the wrong answer ***sorry but those are OS rules, a q per post

OpenStudy (just_one_last_goodbye):

@Somy there really isn't a rule for that

OpenStudy (anonymous):

Can you just turn a blind eye at this, i dont wanna close >_<

OpenStudy (just_one_last_goodbye):

o_o u say that theres a rule for that but... the CoC doesn't say anything

OpenStudy (anonymous):

0-0 so I dont have to close!! okay next question ^_^ Which of the following is the mathematical representation of Gay-Lussac's law? (Points : 3) P1V1 = V2/T2 P1/T1 = P2/T2 V1/T1 = V2/T2 P1V1 = P2V2

OpenStudy (anonymous):

I put B on that

OpenStudy (somy):

Can of soda is also wrong if you chose D

OpenStudy (anonymous):

o-o

OpenStudy (just_one_last_goodbye):

@Somy how so? I know the outer atmosphere doesn't do anything to the can but the atmosphere in the inside of the can yes it does.

OpenStudy (somy):

The pressure of the gas inside the can is lower than that of outside which is why when you open the can the gas particles rush out the question is asking about pressure inside the can so as particles rush out= the pressure will decrease

OpenStudy (somy):

+ @LeilaJudeh is this an online test?

OpenStudy (anonymous):

Nope study guide

OpenStudy (somy):

hmmmm then do you understand each question u've asked so far?

OpenStudy (somy):

why is which answer

OpenStudy (anonymous):

Yes teacher, I have ^_^

OpenStudy (somy):

lol that teacher doe just make sure you learn, & dont cheat it won't take you anywhere in future/adult world T_T

OpenStudy (anonymous):

I know ^_^ Ty for caring <33

OpenStudy (somy):

:)

OpenStudy (somy):

Good luck!

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

so captain you still up to help?

OpenStudy (somy):

sure but please can you open a new q T_T coz its a pain going all the way down

OpenStudy (anonymous):

okay xD